# Effects of Large Optical Depth on CD Hi  21-cm Non-Gaussian Signal

### v1

@@ 3.1 Hi  21-cm optical depth ($\tau$) distribution
We plot the histogram and the cumulative distribution function (CDF) of simulated $\tau$ at redshift $z=12.3$ for three models in Fig. 2. The average $\tau$ values at $z=12.3$ for these models are higher compared to other redshifts, which can be seen in the upper panel of Fig. 1.
Figure 2: This shows histograms of Hi 21-cm optical depth ($\tau$) for models with $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}$ ${\rm M}_{\odot}$ (top left); $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$ (top right) and $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$ (bottom left), at redshift $z=12.3$. The bottom right panel shows the cumulative distribution function (CDF) of $\tau$ for all three scenarios.
We see from the CDF (refer to the bottom right panel of Fig.2) that as many as $\sim 55\%$ of pixels in the simulated cube have $\tau$ value greater than $0.1$ for the model corresponding to $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$. Even for higher X-ray efficiency models corresponding to $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$ and $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$, around $49\%$ and $25\%$ of the total pixels- respectively, where spin temperature is relatively higher- have $\tau$ value greater than $0.1$. Furthermore, we observe that in all scenarios, the distribution of $\tau$ is positively skewed. The optical depth ($\tau$) depends on both the matter density and the spin temperature (refer to equation (2)), both of which are non-Gaussian. This results in a non-Gaussian $\tau$ field, and consequently contributes to a non-Gaussian $\delta T_{b}$ field during CD.
@@ 3.2 Non-Gaussian Hi 21-cm signal
Figure 3: This shows the probability distribution function (PDF) of simulated $\delta T_{\rm b}$ for models with $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}$ ${\rm M}_{\odot}$; $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$ and $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$, at redshift $z=12.3$.
@@ 3.3 Skewness and its redshift evolution
Figure 4: This shows the redshift evolution of skewness for the approximated and exact $\delta T_{\rm b}$ (left panels) and the percentage change (right panels) for all the CD scenarios. In the left panel the blue solid and red dashed lines represent the skewness for the approximated and exact $\delta T_{\rm b}$ respectively.
@@
The left column of Fig.4 shows skewness as a function of redshift $z$ for all scenarios considered here. We can see that the skewness is negative from the beginning of CD until the mean IGM temperature remains lower than the CMBR temperature. Furthermore, the skewness calculated using the accurate $\delta T_{b}$ is smaller in magnitude as compared to the approximated one. The right column of Fig. 4 shows the percentage difference between the two estimates, which quantifies the effect of higher-order terms. We find that all models show significant deviations in skewness. These deviations tend to occur at the redshifts where $\tau$ is large.
At the beginning of CD, skewness is largely negative around redshift $z=18$. As redshift decreases, its magnitude gradually reduces until the heating of the IGM begins. However, during the initial stages of X-ray heating the skewness starts becoming more negative again. As redshift decreases further, heated regions around sources grow significantly and percolates through the IGM. As a result, skewness becomes positive and continues to increase. Eventually, when the entire IGM is nearly heated, the skewness decreases again. For example, in the model with $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$ and $f_{\rm x}=1000$, skewness is $\approx-30$ at redshift $z\approx 18$. Its magnitude gradually decreases until redshift $z\sim 14$ as the $\rm{Ly{\alpha}}$ coupling spreads through the IGM and nears completion. Following this, X-ray heating starts, and skewness tends towards more negative values until redshift $z\approx 12.5$. However, as the heated regions expand and percolates through the IGM, the skewness quickly becomes positive, reaching a peak around $z\approx 11.5$. Later, skewness decreases again as almost the entire IGM gets heated. A similar pattern is observed in other scenarios. However, the actual timings of these processes differ depending on the values of $M_{\rm h,min}$ and $f_{\rm x}$. In general, higher values of $M_{\rm h,min}$ and/or lower values of $f_{\rm x}$ delay these processes.
We can further see from the left column of Fig. 4 that the initial trend in the redshift evolution of skewness is almost similar for a fixed $M_{\rm h,min}$ value when $\rm{Ly{\alpha}}$ coupling dominates. This redshift range is from the beginning up to $z\approx 16.5$ for $M_{\rm h,min}=10^{9}$ ${\rm M}_{\odot}$ and upto $z\sim 14$ for $M_{\rm h,min}=10^{10}$ ${\rm M}_{\odot}$ models. For larger value of $f_{\rm x}$, X-ray heating starts earlier and dominates the $\rm{Ly{\alpha}}$ coupling, shortening its redshift window (lower panel of Fig. 1).
Based on the results above, we conclude that large heated and Ly-$\alpha$ coupled regions enhance the non-Gaussianity of the signal. The completion of $\rm{Ly{\alpha}}$ coupling and IGM heating, however, leads to a gradual decrease in non-Gaussianity.
@@ 3.4 Impact of large optical depth on Skewness
We now shift our focus to the impact of high Hi 21-cm optical depth on $\delta T_{\rm b}$ and its statistics. If we choose a model where heating is inefficient as in $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}$ ${\rm M}_{\odot}$, $\tau$ would be sufficiently large throughout the entire CD as the spin temperature remains very low. As a result, $\delta T_{\rm b}$ calculated using exact equation is considerably suppressed compared to $\delta T_{\rm b}$ calculated using the approximated equation (see lower panel of fig.1). Further, the skewness of $\delta T_{\rm b}$ is suppressed compared to that estimated using the approximated $\delta T_{\rm b}$.
@@
X-ray heating is moderate in the model with $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$. Therefore, the change in the skewness between approximated and exact $\delta T_{\rm b}$ goes up to $\sim 40\%$ at redshift $z\sim 14$ when the average optical depth becomes maximum in the model. Similarly, we see that the change is only up to $\approx 25\%$ at redshift $z\approx 12$ for $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$ for the same X-ray efficiency parameter. Models with $f_{\rm x}=1000$ exhibit similar features, but these appear at relatively earlier redshifts. Here, the maximum difference in skewness between the exact and approximated models is $\approx 60\%$ and $\approx 20\%$ respectively for $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ and $\rm 10^{10}\,{\rm M}_{\odot}$.
@@ 3.5 Redshift evolution of bispectrum
[…]
$k_{1}\geq k_{2}\geq k_{3};\quad{\rm}0.5\leq t,\,\mu\leq 1\quad{\rm and}\quad 2\mu t\geq 1.$
@@
Figure 5: This shows the redshift evolution of the spherically averaged dimensionless bispectrum for approximated and exact $\delta T_{\rm b}$, along with the percentage change between them, at the scale $k_{1}=0.16\,{\rm Mpc}^{-1}$. In the left panel, the blue solid and dashed lines represent bispectrum of approximated and exact $\delta T_{\rm b}$, respectively, for $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ while, the red solid and dashed lines represent the same for $M_{\rm h,min}=10^{10}{\rm M}_{\odot}$. In the right panel, the blue and red solid lines represent the percentage difference between the approximated and exact $\delta T_{\rm b}$ for the models with $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$ and $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$, respectively.
@@
In Figs. 5 and 6, we show the redshift evolution of bispectra for both the exact and approximated $\delta T_{\rm b}$ at two different length scales. Fig. 5 shows results for a large length scale (small $k$) corresponding to $k_{1}=0.16$ Mpc-1 and Fig. 6 shows results for a small length scale corresponding to $k_{1}=0.8$ Mpc-1. We plot the spherically averaged dimensionless bispectrum $\Delta^{3}(k_{1},\mu,t)={k_{1}}^{3}{k_{2}}^{3}B(k_{1},k_{2},k_{3})/(2\pi^{2})^{2}$ (Majumdar et al., 2020) for both the equilateral and squeezed limit triangles. The bispectrum corresponding to the equilateral triangles finds the correlation between equal length scales and that corresponds to the squeezed limit configuration calculates the correlation between two equal length scales and one very large length scale, respectively.
@@
In Fig. 5, we see that $\Delta^{3}(k_{1},\mu,t)$ is initially negative for $f_{\rm x}=0.1$ and $46.4$. At the beginning of CD, the spin temperature $T_{S}$ remains coupled to the CMBR temperature $T_{\gamma}$. This makes the Hi 21-cm signal nearly undetectable. However, as the first sources emerge, $\rm{Ly{\alpha}}$ photons from these sources begin coupling the spin temperature with the IGM kinetic temperature. As IGM kinetic temperature is smaller than the CMBR temperature, $\rm{Ly{\alpha}}$ coupled regions around the first sources appear as regions of negative $\delta T_{\rm b}$ against a zero background signal. This results in a negative bispectrum that continues as long as the $\rm{Ly{\alpha}}$ coupled regions remain isolated. However, these regions expand and percolate through the IGM. Eventually, the entire IGM spin temperature gets coupled to the IGM kinetic temperature. This widespread $\rm{Ly{\alpha}}$ coupling ultimately reduces the non-Gaussianity of the Hi 21-cm signal. We find that the bispectrum is negative (Fig. 5) from the start of CD until the redshifts $z\approx 14$ and $\approx 12$ for $M_{\rm h,min}=10^{9}$ and $10^{10}\,{\rm M}_{\odot}$, respectively, both for the equilateral and squeezed triangle configurations and $f_{\rm x}=0.1$. For a fixed $f_{\rm x}$ value, the overall nature in the features of bispectrum for both halo mass scenarios is almost same. The only difference is that, for higher $M_{\rm h,min}$ value, all the processes start at lower redshift. The redshifts until which the bispectrum remains negative due to $\rm{Ly{\alpha}}$ coupled regions is higher for higher $f_{\rm x}$, as can be seen in Fig. 5. However, we note that the bispectrum is positive since the beginning of CD for $f_{\rm x}=1000$ and the equilateral triangle configuration.
@@
As the $\rm{Ly{\alpha}}$ coupling approaches completion, X-ray radiation from the first sources begins to heat the surrounding IGM. This heating raises both the kinetic temperature and the Hi 21-cm spin temperature of the IGM around these sources above the CMBR temperature. This makes $\delta T_{\rm b}$ positive inside heated regions, although the background IGM remains colder. The presence of large heated regions around the sources, buried in a colder IGM, makes the bispectrum positive. We see, in Fig. 5, that the bispectrum becomes positive at redshifts $z\approx 14$ and $\approx 12$ for $f_{\rm x}=0.1$, corresponding to $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ and $10^{10}{\rm M}_{\odot}$, respectively. The transition redshifts from negative to positive bispectrum are $z\approx 16$ and $\approx 14$ for $f_{\rm x}=46.4$ and $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ and $10^{10}{\rm M}_{\odot}$, respectively, in the equilateral triangle configuration. The transition occurs slightly later, at $z\approx 15$ and $\approx 13$, for the squeezed triangle configuration with the same halo mass thresholds. For $f_{\rm x}=1000$, X-ray heating is highly efficient, and the bispectrum remains positive from the beginning of CD in the equilateral triangle configuration. However for the squeezed triangle configuration, the bispectrum is initially negative for a short period of time, and becomes positive as X-ray heating starts. We see in this study that, for squeezed limit configuration, all the features appear at later stages compared to the equilateral one. The reason is that, in the squeezed-limit triangle configuration, there is a correlation involving one very large length scale, and astrophysical processes begin to affect such large scales at later times, i.e., at lower redshifts.
@@
As the X-ray heating progresses, it gradually percolates through the IGM. This raises the entire IGM temperature above the CMBR temperature. The widespread heating again reduces the bispectrum. We see this effect in Fig. 5 for $f_{\rm x}=46.4$, for both equilateral and squeezed triangle configurations. Toward the final stage of heating, almost the entire IGM becomes heated, except a few underdense regions that are located far from the sources. This creates a scenario where $\delta T_{\rm b}$ is positive in most regions of the IGM but remains negative in those cold and underdense isolated pockets. This results in a large-scale negative bispectrum. For example, for $f_{\rm x}=46.4$ and $1000$, we see that the bispectrum becomes negative again at later times.

### v2

@@ 3.1 Hi  21-cm optical depth ($\tau$) distribution
We plot the probability distribution function (PDF) and the cumulative distribution function (CDF) of simulated $\tau$ at redshift $z=12.3$ for three models in Fig. 2. The average $\tau$ values at $z=12.3$ for these models are higher compared to other redshifts, which can be seen in the upper panel of Fig. 1.
Figure 2: Probability distribution function (PDF) of Hi 21-cm optical depth ($\tau$) for models with $f_{\rm x}=0.1$, $_{\rm h,min}=10^{9} M{\rm M}_{\odot}$ (top left); $f_{\rm x}=0.1$, $_{\rm h,min}=10^{10} M{\rm M}_{\odot}$ (top right) and $f_{\rm x}=46.4$, $_{\rm h,min}=10^{10} M{\rm M}_{\odot}$ (bottom left), at redshift $z=12.3$. The bottom right panel shows the cumulative distribution function (CDF) of $\tau$ for all three scenarios.
We see from the CDF (refer to the bottom right panel of Fig.2) that as many as $\sim 55\%$ of pixels in the simulated cube have $\tau$ value greater than $0.1$ for the model corresponding to $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$. For models with $f_{\rm x}=0.1$, $M_{\rm h,min}=10^{10},{\rm M}_{\odot}$ and $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{10},{\rm M}_{\odot}$, about $49\%$ and $25\%$ of the total pixels, respectively, have $\tau>0.1$. Furthermore, we observe that in all scenarios, the distribution of $\tau$ is positively skewed. The optical depth ($\tau$) depends on both the matter density and the spin temperature (refer to equation (2)), both of which are non-Gaussian. This results in a non-Gaussian $\tau$ field, and consequently contributes to a non-Gaussian $\delta T_{b}$ field during CD.
@@ 3.2 Non-Gaussian Hi 21-cm signal
Figure 3: Probability distribution function (PDF) of simulated $\delta T_{\rm b}$ for models with $f_{\rm x}=0.1$, $_{\rm h,min}=10^{9} M{\rm M}_{\odot}$; $f_{\rm x}=0.1$, $_{\rm h,min}=10^{10} M{\rm M}_{\odot}$ and $f_{\rm x}=46.4$, $_{\rm h,min}=10^{10} M{\rm M}_{\odot}$, at redshift $z=12.3$.
@@ 3.3 Skewness and its redshift evolution
Figure 4: Redshift evolution of skewness for the approximated and exact $\delta T_{\rm b}$ (left panels) and the percentage change (right panels) for all the CD scenarios. In the left panel the blue solid and red dashed lines represent the skewness for the approximated and exact $\delta T_{\rm b}$ respectively.
@@
The left column of Fig. 4 shows skewness as a function of redshift $z$ for all scenarios considered here. We can see that at the beginning of CD, skewness is largely negative around redshift $z=18$. As redshift decreases, its magnitude gradually reduces until the heating of the IGM begins. However, during the initial stages of X-ray heating the skewness starts becoming more negative again. As redshift decreases further, heated regions around sources grow significantly and percolates through the IGM. As a result, skewness becomes positive and continues to increase. Eventually, when the entire IGM is nearly heated, the skewness decreases again. For example, in the model with $_{\rm h,min}=10^{10} M{\rm M}_{\odot}$ and $f_{\rm x}=1000$, skewness is $\approx-30$ at redshift $z\approx 18$. Its magnitude gradually decreases until redshift $z\sim 14$ as the $\rm{Ly{\alpha}}$ coupling spreads through the IGM and nears completion. Following this, X-ray heating starts, and skewness tends towards more negative values until redshift $z\approx 12.5$. However, as the heated regions expand and percolates through the IGM, the skewness quickly becomes positive, reaching a peak around $z\approx 11.5$. Later, skewness decreases again as almost the entire IGM gets heated. A similar pattern is observed in other scenarios. However, the actual timings of these processes differ depending on the values of $M_{\rm h,min}$ and $f_{\rm x}$. In general, higher values of $M_{\rm h,min}$ and/or lower values of $f_{\rm x}$ delay these processes.
We can further see from the left column of Fig. 4 that the initial trend in the redshift evolution of skewness is almost similar for a fixed $M_{\rm h,min}$ value when $\rm{Ly{\alpha}}$ coupling dominates. This redshift range is from the beginning up to $z\approx 16.5$ for $_{\rm h,min}=10^{9} M{\rm M}_{\odot}$ and upto $z\sim 14$ for $_{\rm h,min}=10^{10} M{\rm M}_{\odot}$ models. For larger value of $f_{\rm x}$, X-ray heating starts earlier and dominates the $\rm{Ly{\alpha}}$ coupling, shortening its redshift window (lower panel of Fig. 1).
@@ 3.4 Impact of large optical depth on Skewness
We now shift our focus to the impact of high Hi 21-cm optical depth on $\delta T_{\rm b}$ and its statistics. If we choose a model where heating is inefficient as in $f_{\rm x}=0.1$, $_{\rm h,min}=10^{9} M{\rm M}_{\odot}$, $\tau$ would be sufficiently large throughout the entire CD as the spin temperature remains very low. As a result, $\delta T_{\rm b}$ calculated using exact equation is considerably suppressed compared to $\delta T_{\rm b}$ calculated using the approximated equation (see lower panel of fig.1). Further, the skewness of $\delta T_{\rm b}$ is suppressed compared to that estimated using the approximated $\delta T_{\rm b}$.
@@
X-ray heating is moderate in the model with $f_{\rm x}=46.4$, $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$. Therefore, the change in the skewness between approximated and exact $\delta T_{\rm b}$ goes up to $\sim 40\%$ at redshift $z\sim 14$ when the average optical depth becomes maximum in the model. Similarly, we see that the change is only up to $\approx 25\%$ at redshift $z\approx 12$ for $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$ for the same X-ray efficiency parameter. Models with $f_{\rm x}=1000$ exhibit similar features, but these appear at relatively earlier redshifts. Here, the maximum difference in skewness between the exact and approximated models is $\approx 60\%$ and $\approx 20\%$ respectively for $(f_{\rm x},M_{\rm h,min})=(1000,10^{9}\,{\rm M}_{\odot})$ and $(1000,10^{10}\,{\rm M}_{\odot})$.
@@ 3.5 Redshift evolution of bispectrum
where $\Delta_{\rm b}(\textbf{{k}})$ is the Fourier conjugate of $\delta T_{\rm b}(\textbf{{x}})$, $\langle\cdots\rangle$ denotes an ensemble average. The Kronecker delta function $\delta^{\rm K}_{3\rm D}$ in the definition is an implication of statistically homogeneity of the signal which ensures that k-triplets must form a closed triangle for a non zero bispectrum. We computed bispectrum of the simulated $\delta T_{\rm b}$ maps using an FFT-based fast estimator (Shaw et al., 2021) where the Kronecker delta function is replaced by summation of plane waves over grid points in real space. To determine the shape of the triangle formed by the three k-modes, this algorithm uses the parametrization: $\mu=-(\textbf{{k}}_{1}\cdot\textbf{{k}}_{2})/(k_{1}~k_{2})$ and $t=k_{2}/k_{1}$, where $k_{i}=|\textbf{{k}}_{i}|$. The cosine of the angle between $\textbf{{k}}_{1}$ and $\textbf{{k}}_{2}$ is denoted as $\mu$, whereas $t$ is the ratio between them and each set of $(k_{1},\mu,t)$ represents a different triangle configuration. For unique triangle configuration, order of three $k$ sides and the range of $\mu$ and $t$ values are prescribed in (Bharadwaj et al., 2020) as,
$k_{1}\geq k_{2}\geq k_{3};\quad{\rm}0.5\leq t,\,\mu\leq 1\quad{\rm and}\quad 2\mu t\geq 1.\$
@@
Figure 5: Redshift evolution of the spherically averaged dimensionless bispectrum for approximated and exact $\delta T_{\rm b}$, along with the percentage change between them, at the scale $k_{1}=0.16\,{\rm Mpc}^{-1}$. In the left panel, the blue solid and dashed lines represent bispectrum of approximated and exact $\delta T_{\rm b}$, respectively, for $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ while, the red solid and dashed lines represent the same for $M_{\rm h,min}=10^{10}{\rm M}_{\odot}$. In the right panel, the blue and red solid lines represent the percentage difference between the approximated and exact $\delta T_{\rm b}$ for the models with $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$ and $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$, respectively.
@@
In Figs. 5 and 6, we show the redshift evolution of bispectra for both the exact and approximated $\delta T_{\rm b}$ at two different length scales. Fig. 5 shows results for a large length scale (small $k$) corresponding to $k_{1}=0.16$ Mpc-1 and Fig. 6 shows results for a small length scale corresponding to $k_{1}=0.8$ Mpc-1. We plot the spherically averaged dimensionless bispectrum $\Delta^{3}(k_{1},\mu,t)={k_{1}}^{3}{k_{2}}^{3}B(k_{1},k_{2},k_{3})/(2\pi^{2})^{2}$ (Majumdar et al., 2020) for both the equilateral and squeezed limit triangles. The bispectrum corresponding to the equilateral triangles finds the correlation between three equal length scales. In contrast, the bispectrum corresponds to the squeezed limit configuration calculates the correlation between two equal length scales and one very large length scale.
@@
In Fig. 5, we see that $\Delta^{3}(k_{1},\mu,t)$ is initially negative for $f_{\rm x}=0.1$ and $46.4$. At the beginning of CD, the spin temperature $T_{S}$ remains coupled to the CMBR temperature $T_{\gamma}$. This makes the Hi 21-cm signal nearly undetectable. However, as the first sources emerge, $\rm{Ly{\alpha}}$ photons start emitting from these sources as a result, $\rm{Ly{\alpha}}$ coupled regions from these sources appear as regions of negative $\delta T_{\rm b}$ against a zero background signal. This results in a negative bispectrum that continues as long as the $\rm{Ly{\alpha}}$ coupled regions remain isolated. However, later these regions expand and percolate through the IGM. This widespread $\rm{Ly{\alpha}}$ coupling ultimately reduces the non-Gaussianity of the Hi 21-cm signal. For $f_{\rm x}=0.1$, we find that the bispectrum is negative (Fig. 5) from the start of CD until the redshifts $z\approx 14$ and $\approx 12$ for $M_{\rm h,min}=10^{9}$ and $10^{10}\,{\rm M}_{\odot}$, respectively, both for the equilateral and squeezed triangle configurations. For a fixed $f_{\rm x}$ value, the overall nature in the features of bispectrum for both halo mass scenarios is almost same. The only difference is that, for higher $M_{\rm h,min}$ value, all the processes start at lower redshift. For higher values of $f_{x}$, bispectrum turns positive earlier, as can be seen in Fig. 5. However, we note that the bispectrum is positive since the beginning of CD for $f_{\rm x}=1000$ and the equilateral triangle configuration.
@@
As the $\rm{Ly{\alpha}}$ coupling approaches completion, X-ray radiation from the first sources begins to heat the surrounding IGM. The presence of large heated regions around the sources, buried in a colder IGM, makes the bispectrum positive. We see, in Fig. 5, that the bispectrum becomes positive at redshifts $z\approx 14$ and $\approx 12$ for $f_{\rm x}=0.1$, corresponding to $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ and $10^{10}{\rm M}_{\odot}$, respectively. The transition redshifts from negative to positive bispectrum are $z\approx 16$ and $\approx 14$ for $f_{\rm x}=46.4$ and $M_{\rm h,min}=10^{9}{\rm M}_{\odot}$ and $10^{10}{\rm M}_{\odot}$, respectively, in the equilateral triangle configuration. The transition occurs slightly later, at $z\approx 15$ and $\approx 13$, for the squeezed triangle configuration with the same halo mass thresholds. For $f_{\rm x}=1000$, X-ray heating is highly efficient, and the bispectrum remains positive from the beginning of CD in the equilateral triangle configuration. However for the squeezed triangle configuration, the bispectrum is initially negative for a short period of time, and becomes positive as X-ray heating starts. We see in this study that, for squeezed limit configuration, all the features appear at later stages compared to the equilateral one. The reason is that, in the squeezed-limit triangle configuration, there is a correlation involving one very large length scale, and astrophysical processes begin to affect such large scales at later times, i.e., at lower redshifts.
@@
As the X-ray heating progresses, it gradually percolates through the IGM. The widespread heating again reduces the bispectrum. We see this effect in Fig. 5 for $f_{\rm x}=46.4$, for both equilateral and squeezed triangle configurations. Toward the final stage of heating, almost the entire IGM becomes heated, except a few underdense regions that are located far from the sources. This creates a scenario where $\delta T_{\rm b}$ is positive in most regions of the IGM but remains negative in those cold and underdense isolated pockets. This results in a large-scale negative bispectrum. For example, for $f_{\rm x}=46.4$ and $1000$, we see that the bispectrum becomes negative again at later times.